Use the following results from a test for marijuana​ use, which is provided by a certain drug testing company. Among 146 subjects with positive test​ results, there are 23 false positive results. Among 151 negative​ results, there are 3 false negative results. Complete parts​ (a) through​ (c). (Hint: Construct a​ table.) a. How many subjects were included in the​ study? The total number of subjects in the study was 297. b. How many subjects did not use​ marijuana? A total of nothing subjects did not use marijuana.
Answer #1

b) Number of subjects who did not used marijuana = 23 + (151 - 3) = 23 + 148 = 171